what is the half life of an isotope that decays to 12.5% of its original activity in 33.6 hours
Answer #1

K   = 2.303/t log[A0]/[A]

t = 33.6 hours

[A0]   = 100

[A]    = 12.5

K    = 2.303/33.6 log(100/12.5)

         = 2.303*0.9030/33.6 = 0.0619 hour^-1

t1/2 = 0.693/K

          = 0.693/0.0619 = 11.2hour

second method

(1/2)^n    = 0.125

(0.5)^n    = (0.5)^3

n     = 3

n   = total time/half life time

3     = 33.6/ half life time

half life time = 33.6/3   = 11.2 hours >>>>answer